OpenStudy (anonymous):

The shorter lef of right triangle is 17cm less than the longer leg. If the hypotenuse is 25cm long, find the perimeter of the triangle

OpenStudy (anonymous):

a = b-17 and c=25 (b-17)^2 + b^2 = c^2 = 625 put into standard quadratic equation form =0, then find b then perimeter is a + b + c.